Hydrofluoric acid, HF, has a Ka of 6.8 ×10−4. What are [H3O+], [F], and [OH] in 0.880 M HF?
Answer #1

Answer

[H3O+] = 0.02412 M

[F-] = 0.02412 M

[OH-] = 4.15 ×10-13 M

Explanation

Dissociation equilibrium of HF is

HF(aq) + H2O(aq) <--------> H3O+(aq) + F-(aq)

Ka = [H3O+][F-] /[HF] = 6.8 ×10-4

Initial concentration

[HF] = 0.880

[H3O+] = 0

[F-​​​​​​] = 0

Change in concentration

[HF] = - x

[H3O+] = +x

[F-] = + x

Equilibrium concentration

[HF] = 0.880 - x

[H3O+] = x

[F-] = x

so,

x2/(0.880 - x) = 6.8 ×10-4

solving for x

x = 0.02412

[HA] = 0.880 - 0.02412= 0.8559 M

[F-] = 0.02412 M

[H3O+] = 0.02412 M

[H3O+] [OH-] = Kw

Kw is ionic product of water , 1.00 × 10-14

0.02412 M × [OH-] = 1.00 ×10-14M2

[OH-] = 4.15 ×10-13M